OpenStudy (anonymous):

Which number should replace the question mark? 13 8 3 7 4 1 5 1 8 10 9 2 46 32 26 3 14 ? 11 3

OpenStudy (anonymous):

33-14=19

OpenStudy (anonymous):

check those 13+8=3*7 4+1=5*1 8+10=9*2 46+32=26*3